OpenStudy (styxer):

When x+y = 120º , x>= 0º and y>=0º , the maximum of sin x + sin y is [??] and the minimum of that is [???] (Just to make sure that I did the right way, please)

OpenStudy (styxer):

So, I considerate the following values : 120,90,60,30,0 sin 120 = sqrt(3)/2 sin 90 = 1 sin 60 = sqrt(3)/2 sin 30 = 1/2 sin 0 = 0 And just combined the values to find the maximum (which is sin 60 + sin 60) and the minimum (which is sin 120 + sin 0) Is this enough to guarantee both minimum and maximum?

ganeshie8 (ganeshie8):

You're considering only a few combinations

ganeshie8 (ganeshie8):

familiar with the identity \[a\cos x+b\sin x = \sqrt{a^2+b^2}\sin(\star)\] ?

OpenStudy (styxer):

No... :(

ganeshie8 (ganeshie8):

How about calculus ?

OpenStudy (styxer):

I know the basics... Which part?

ganeshie8 (ganeshie8):

Familiar with finding the max/min of a function by setting the first derivative equal to 0 ?

OpenStudy (styxer):

Yes

ganeshie8 (ganeshie8):

You want to find the min/max of sin(x) + sin(120-x) over the interval [0, 120]

ganeshie8 (ganeshie8):

Let f(x) = sin(x) + sin(120-x) f'(x) = ?

OpenStudy (styxer):

f'(x) = cos(x) + cos(120-x)

ganeshie8 (ganeshie8):

Careful, derivative of sin(120-x) is not just cos(120-x)

ganeshie8 (ganeshie8):

Remember the chain rule ?

ganeshie8 (ganeshie8):

You must multiply by the derivative of `120-x`

ganeshie8 (ganeshie8):

What's the derivative of `120-x` with respect to `x` ?

OpenStudy (styxer):

it's -1 ?

ganeshie8 (ganeshie8):

Yes f(x) = sin(x) + sin(120-x) f'(x) = ?

OpenStudy (styxer):

f'(x) = cos(x) - cos(120-x)

ganeshie8 (ganeshie8):

Yep, set that equal to 0 and solve x over the given interval

ganeshie8 (ganeshie8):

cos(x) - cos(120-x) = 0 cos(x) = cos(120-x)

ganeshie8 (ganeshie8):

Recall that cos(x) = cos(-x)

ganeshie8 (ganeshie8):

cos(x) - cos(120-x) = 0 cos(x) = cos(120-x) This is same as cos(-x) = cos(120-x) -x = 120-x x = ?

OpenStudy (styxer):

x = 120

ganeshie8 (ganeshie8):

really ?

ganeshie8 (ganeshie8):

try again

OpenStudy (styxer):

x=60

ganeshie8 (ganeshie8):

Yes, that means x = 60 is a critical value. Since [0, 120] is the range, other critical values are x=0, x=120. Calculus guarantees us that the max/min values of f(x) can occur only at these critical values.

ganeshie8 (ganeshie8):

So it is sufficient to check the function at these 3 critical values : x = 0, 60, 120

OpenStudy (styxer):

And then I just need to do sin(x) + sin(120-x) with these three values, right?

ganeshie8 (ganeshie8):

Yes

OpenStudy (styxer):

I didn't think about y being 120-x , that's interesting

OpenStudy (styxer):

Thank you!! :)

ganeshie8 (ganeshie8):

Np :) But I prefer solving this problem using the identity \[a\cos x+b\sin x = \sqrt{a^2+b^2}\sin(\star)\] This is my favorite identity from trig. You will see it a lot if you take differential equations...

ganeshie8 (ganeshie8):

Look it up when free http://www.mash.dept.shef.ac.uk/Resources/web-rcostheta-alphaetc.pdf

OpenStudy (styxer):

I will, thanks